OpenStudy (anonymous):

The distance between Chicago and Cleveland is 354 miles. If a person leaves Chicago at 9:50 am Central Time and arrives in Cleveland at 5:30pm the same day Eastern Time, at what average speed does he travel, correct to the nearest mile? (Central time is one hour earlier than Easter Time. (A) 46mph (B) 50mph (C) 53mph (D) 55mph (E) 56mph

OpenStudy (anonymous):

What answer do you think it is?

OpenStudy (anonymous):

I think it's...A..

OpenStudy (anonymous):

It is C

OpenStudy (anonymous):

How do you get C?

OpenStudy (anonymous):

Turn 6 hour and 40 minutes into decimals and it = 6.667, which is your time

OpenStudy (anonymous):

Pretty much you are using the formula d= rt

OpenStudy (anonymous):

You are trying to find the speed/rate which is r

OpenStudy (anonymous):

you have distance = d and you have the time = t

OpenStudy (anonymous):

To find the rate just d = rt d/t = r r = d/t r = 354/6.667 which gives you your answer

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

yeah. Thanks so much!! ^^

OpenStudy (anonymous):

\[r = \frac{d}{t} \]\[r = \frac{354}{17.5-\left(9+\frac{50}{60}\right)}= 46.17=46 \text{ mph} \text{ rounded} \]

OpenStudy (anonymous):

17.5 is 5:30 PM

OpenStudy (anonymous):

robtobey the time is only 6.667. Your calculation is wrong.

OpenStudy (anonymous):

Reread the problem. It is going from central to est

OpenStudy (anonymous):

@MDoodler You are correct. Rushed through the solution and missed the time zone corrections.